OpenStudy (anonymous):

Find the domain, range and vertex of the following parabola. y=x^2-14x+44 Show all work.

OpenStudy (dtan5457):

First you need to put that standard form quadratic to vertex form

OpenStudy (dtan5457):

Know how to do that?

OpenStudy (anonymous):

Not really, sorry.

OpenStudy (dtan5457):

Ok. standard form was x^2-14x+44 you need to complete the square. start by subtracting the value of c (44) y-44=x^2-14x now, complete the perfect square trinomial. use the formula (b/2)^2 (-7)^2=49 add that to both sides y+5=x^2-14x+49 simplify to y+5=(x-7)^2 y=(x-7)^2

OpenStudy (dtan5457):

y=(x-7)^2-5 to get the vertex from this, get the number from inside the parenthesis, switch the sign, and -5 will be the y value. vertex: 7,-5

OpenStudy (dtan5457):

got that?

OpenStudy (dtan5457):

im still here :)